# Stellar/BH Population in AGN Disks: Direct Binary Formation from Capture Objects in Nuclei Clusters

### v3

@@ 2.1.1 Inclination damping and eccentricity evolution
Figure 3: $\dot{e}$ as a function of inclination for different $e$. The upper panel shows the case with $\cos\omega=0$ where the two crossing points have the same distance to the SMBH. The bottom panel shows the case with $\cos\omega=\pm 1$ where the periapsis and apoapsis lie within the disk. Eccentricity will always be excited for retrograde orbits with high inclination. There’s no parameter space for eccentricity damping with small $\cos\omega$. As $\cos\omega$ increases, low inclination prograde orbits could undergo eccentricity damping.
@@ 2.1.3 Integral of motion
The general picture for the stellar capture is shown in Figure 5, where for a given semi-major axis, three critical angles $I_{c}$, $I_{\rm TDE}$ and $I_{\rm e}$ divide the parameter space into three different regimes. For very high inclination orbits, the semi-major axis shrinks very fast by a factor of $\frac{1}{(1-e_{0}^{2})\cos^{4}(I_{0}/2)}$. This contributes to the TDE by disk captures. For retrograde orbits not associated with extremely high inclination, their semi-major axis damping would be much slower and the eccentricity excitation becomes relevant until the inclination decreases to the critical angle $I_{\rm e}$. Then the eccentricity starts to decrease fast and eventually becomes nearly zero as the stars are captured by the disk.
@@ 2.2.1 Capture, eccentricity damping, and excitation
Figure 7: Similar to Figure 3, but for gas dynamical friction. Unlike the aerodynamical drag, there is more parameter space for eccentricity excitation, there is more parameter space for eccentricity damping in the gas dynamical friction case. The vertical dashed lines show the critical inclination for eccentricity excitation/damping indicated by Equation 69.
Figure 7 illustrates the $\dot{e}$ as a function of inclination for different $e$ and $\omega$. Similar to Figure 3, the upper panel/bottom panel shows the case for $\cos\omega=0$ and ±$\pm 1$, respectively. Different from the aerodynamical drag, for $\cos\omega=0$, the eccentricity always decay, even if the orbit is purely retrograde. As $\cos I_{\rm e}$ becomes larger than -1, retrograde orbits with inclination $\cos I<\cos I_{\rm e}$ will undergo eccentricity excitation. For eccentricity damping to appear, we need $\cos I_{\rm e}>-1$, such that any inclination $I<I_{\rm e}$ will undergo eccentricity damping. The criterion for $\cos I_{\rm e}>-1$ is $|\cos\omega|\sim>0.1$ for various eccentricities. Therefore, there’s only a small region in the parameter space where eccentricity damping never appears.
@@ 3.1.1 $\alpha$-disk
The $\alpha$ disk model can be characterized by three parameters: the accretion rate efficiency ($\lambda_{\rm d}$), the Toomre parameter ($Q_{\rm d}$), and the viscosity parameter ($\alpha_{\rm d}$), which drives the disk instability. The accretion rate can be approximated as follows:
@@
$v_{r}$ $=$ $\alpha h^{2}\sqrt{\frac{GM_{\rm SMBH}}{r}}$ (78)
@@ 4.2 Mass filling function of captured objects
*$\mathcal{M}_{*}(t)$ $\sim$ $2M_{\rm SMBH}f_{*}\bigg{(}\frac{\Sigma_{m}}{\Sigma_{*}}\frac{t}{T_{m}}\bigg{)}^{1-\gamma_{\rm NSC}/3}$ (86)
@@
Figure 17 shows the captured total mass as a function of time for different disk models and NSC density profiles from our MC simulations. The upper panel shows the cases for stars with $f_{*}=1$ and the bottom panel shows the cases for BHs with $f_{\rm BH}=1$. As indicated, the mass filling function for stars *$\mathcal{M}_{*}\propto t^{1-\gamma_{\rm NSC}}$ and the mass filling function for BHs $\mathcal{M}_{\rm BH}\propto t$.
@@ 5 Conclusions
  • If the AGN disk is a self-gravitating disk with a surface density scaling as $\Sigma\propto r^{-3/2}$, besides the initially embedded stars, the total mass of the captured stars can be well described by a mass filling function *$\mathcal{M}_{*}(t)=2M_{\rm SMBH}f_{*}\left(\frac{\Sigma_{m}}{\Sigma_{*}}\frac{t}{T_{m}}\right)^{1-\gamma_{\rm NSC}/3}$. The captured stars will be distributed between $r_{g}$ and $r_{m}\left(\frac{\Sigma_{m}}{\Sigma_{*}}\frac{t}{T_{m}}\right)^{1/3}$, with a number profile $\mathrm{d}N/\mathrm{d}r\propto r^{-1/4}$.

### v5

@@ 2.1.1 Inclination damping and eccentricity evolution
Figure 3: $\dot{e}$ as a function of inclination for different $e$. The upper panel shows the case with $\cos\omega=0$ where the two crossing points have the same distance to the SMBH. The bottom panel shows the case with $\cos\omega=\pm 1$ where the periapsis and apoapsis lie within the disk. Eccentricity will always decrease for all inclinations.
@@ 2.1.3 Integral of motion
The general picture for the stellar capture is shown in Figure 5, where for a given semi-major axis, two critical angles $I_{c}$ and $I_{\rm TDE}$ divide the parameter space into two different regimes. For very high inclination orbits, the semi-major axis shrinks very fast by a factor of $\frac{1}{(1-e_{0}^{2})\cos^{4}(I_{0}/2)}$. This contributes to the TDE by disk captures. For retrograde orbits not associated with extremely high inclination, their semi-major axis damping would be slower and associated with eccentricity decrease.
@@ 2.2.1 Capture, eccentricity damping, and excitation
Figure 7: Similar to Figure 3, but for gas dynamical friction. Unlike the aerodynamical drag, there is parameter space for eccentricity excitation. The vertical dashed lines show the critical inclination for eccentricity excitation/damping indicated by Equation 69.
Figure 7 illustrates the $\dot{e}$ as a function of inclination for different $e$ and $\omega$. Similar to Figure 3, the upper panel/bottom panel shows the case for $\cos\omega=0$ and ±$\pm 1$, respectively. For $\cos\omega=0$, similar to the aerodynamical drag, the eccentricity always decays, even if the orbit is purely retrograde. As $\cos I_{\rm e}$ becomes larger than -1, retrograde orbits with inclination $\cos I<\cos I_{\rm e}$ will undergo eccentricity excitation. For eccentricity damping to appear, we need $\cos I_{\rm e}>-1$, such that any inclination $I<I_{\rm e}$ will undergo eccentricity damping. The criterion for $\cos I_{\rm e}>-1$ is $|\cos\omega|\sim>0.1$ for various eccentricities. Therefore, there’s only a small region in the parameter space where eccentricity damping never appears.
@@ 3.1.1 $\alpha$-disk
The $\alpha$ disk model can be characterized by three parameters: the accretion rate efficiency ($\lambda_{\rm d}$), the Toomre parameter ($Q_{\rm d}$), and the viscosity parameter ($\alpha_{\rm d}$). The accretion rate can be approximated as follows:
@@
$v_{r}$ $=$ $\alpha_{d}h^{2}\sqrt{\frac{GM_{\rm SMBH}}{r}}$ (78)
@@ 4.2 Mass filling function of captured objects
$\mathcal{M}_{*}(t)$ $\sim$ $2M_{\rm SMBH}f_{*}\bigg{(}\frac{\Sigma_{m}}{\Sigma_{*}}\frac{t}{T_{m}}\bigg{)}^{1-\gamma_{\rm NSC}/3}$ (86)
@@
Figure 17 shows the captured total mass as a function of time for different disk models and NSC density profiles from our MC simulations. The upper panel shows the cases for stars with $f_{*}=1$ and the bottom panel shows the cases for BHs with $f_{\rm BH}=1$. As indicated, the mass filling function for stars $\mathcal{M}_{*}\propto t^{1-\gamma_{\rm NSC}}$ and the mass filling function for BHs $\mathcal{M}_{\rm BH}\propto t$.
@@ 5 Conclusions
  • If the AGN disk is a self-gravitating disk with a surface density scaling as $\Sigma\propto r^{-3/2}$, besides the initially embedded stars, the total mass of the captured stars can be well described by a mass filling function $\mathcal{M}_{*}(t)=2M_{\rm SMBH}f_{*}\left(\frac{\Sigma_{m}}{\Sigma_{*}}\frac{t}{T_{m}}\right)^{1-\gamma_{\rm NSC}/3}$. The captured stars will be distributed between $r_{g}$ and $r_{m}\left(\frac{\Sigma_{m}}{\Sigma_{*}}\frac{t}{T_{m}}\right)^{1/3}$, with a number profile $\mathrm{d}N/\mathrm{d}r\propto r^{-1/4}$.